OpenStudy (idealist10):

tan^-1 y=tan^-1 x+C Solve for y.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

No initial values? Anyway, just take the tangent of both sides: \[\tan(\tan^{-1}y)=\tan(\tan^{-1}x+C)~~\iff~~y=\tan(\tan^{-1}x+C)\]

OpenStudy (idealist10):

Wait a minute. Let me work it out.

OpenStudy (idealist10):

This is the full problem: Solve y'=(1+y^2)/(1+x^2) explicitly. HINT: Use the identity tan(A+B)=(tanA+tanB)/(1-tanAtanB).

OpenStudy (anonymous):

Okay, so using the identity gives you \[\begin{align*} y&=\tan(\tan^{-1}x+C)\\\\ &=\frac{\tan(\tan^{-1}x)+\tan C}{1-\tan(\tan^{-1}x)\tan C}\\\\ &=\frac{x+\tan C}{1-x\tan C}\\\\ \end{align*}\]

OpenStudy (idealist10):

Thank you so much! Sorry the internet runs a bit slow.

OpenStudy (anonymous):

You're welcome!